OpenStudy (anonymous):

Solve C = 3(2f – 5h) for f in terms of C and h. You must show all work. ASAP!

OpenStudy (mertsj):

C/3=2f-5h C/3+5h=2f C/6+5/2h=f

OpenStudy (anonymous):

Distribute, add 15h to both sides, divide by six to solve for f in terms of the other variables.

OpenStudy (anonymous):

...or do it the way Mertsj did....

OpenStudy (anonymous):

C = 3(2f – 5h) C=6f-15h +15h +15h 15h+C=6f 6f=(15h+C) /6 /6 f=(15h+C)/6 or 5h/2 +C/6